# Hawking radiation in a charged, cosmological black hole

### v1

@@ III.2 Effective temperature perceived by a radial freefaller with positive energy
Before proceeding to examine the extremal conditions of the RNdS metric, we may understand the features of the effective temperature in a Schwarzschild de Sitter (SdS) model, which is a simplification (Q=0) of the same metric. It has only the event and cosmological horizons, thereby eliminating the intricacies that accompany the Cauchy horizon. It is therefore a good way to examine the individual effects of the mass and cosmological expansion on the effective temperature. The first graph, on the upper left of Figure 1, shows the effective temperature for an SdS black hole as a function of the observer’s position. Both the ingoing and outgoing modes are continuous across the event and cosmological horizons, in accordance with the Eqs. (25) and (26). This means that an ingoing(outgoing) observer will be able to cross the event(cosmological) horizon without noticing a discontinuity in the effective temperature. Neither the event nor the cosmological horizons act as a physical barrier, and are simply the radii at which space flows inwards (or outwards) at the speed of light. As the observer moves to a predominantly de Sitter (dS) space, the effective temperatures of the ingoing and outgoing modes converge to $1/L_{dS}$. In a purely dS space, then, the observer sees an isotropic equilibrium in both the modes, with the value of the effective temperature in this regime being independent of both the (positive) energy and the position of the observer. The energy-independence is in support of the observer-dependence of particle production in dS space [18]. The position-independence reinforces a homogeneity in the dS spacetime. At the other extreme, as the observer approaches the singularity, they will see a monotonic increase, and an asymptotic convergence in the positive effective temperatures of the modes. In the effective temperature formalism, a positive effective temperature corresponds to a perceived positive rate of redshift in the modes. As the curvature of spacetime increases towards the singularity, both the modes are pulled inwards, and an infalling observer perceives an increasing rate of redshift. So also, the redshift in the modes in a purely dS spacetime may be understood as resulting from the increasing rate of the expansion of space in this regime. The mass of the black hole and the cosmological expansion, individually, only cause a redshift. The blueshifting, that is seen in the other graphs of Figure 1, must then be attributed entirely to the effects of the inner horizon.
@@
A very typical representation of an RNdS black hole is seen in the upper right graph of Figure 1, with parametric values of $M=0.05$, $Q=0.02$. Away from the inner horizon, it is fairly similar to the SdS case. Once again, there is a continuity across the event and cosmological horiozns, and the modes converge to unity in asymptotically dS space. The key feature of this graph, however, comes from the negative effective temperatures of the modes perceived by an observer approaching the inner horizon. A similarity in the cosmological and inner horizons stems from the fact that both have negative surface gravities. But unlike at the cosmological horizon, an observer who started from rest at the maximum of the horizon function and is subject to no external force agents, near the inner horizon is infalling (and ingoing). Effectively, the observer is falling inwards and the modes are being repelled by the surface, resulting in a blueshift (which translates to a negative effective temperature). The transition from positive to negative effective temperatures must occur when the redshift caused by the event and cosmological horizons is overcome by the blueshift due to the inner horizon. This in turn, should have a dependency on the closeness of $r_{+}$ and $r_{-}$. Mathematically, the transition occurs at $\varkappa_{ob}=\varkappa_{+}$ for the outgoing modes and $\varkappa_{ob}=\varkappa_{c}$ for the ingoing modes. While the negative surface gravity certainly causes a blueshift, the effects are much stronger on the outgoing modes than on the ingoing modes because the repulsion supplements their intended direction. As a result, the outgoing modes diverge negatively, while the ingoing modes are blueshifted, but remain finite across the inner horizon.
@@
Having studied the general properties of a very typical RNdS black hole, we may now turn to the more extremal situations. A sharkfin diagram representing the physically permitted parameter space for the metric is plotted in Figure 2. There are primarily, three regions of interest: the SdS limit (Q=0), the charged Nariai limit ($r_{+}=r_{c}$) and the super-extremal limit ($r_{+}=r_{-}$). In asymptotically flat spacetime, setting M=Q results in an extremal black hole, with a degeneracy in the event and inner horizons. The existence of the cosmological constant raises the upper bounds of the extremality, so that black holes may exist in a limited range where $Q>M$ (the cold region). M=Q (referred to as the lukewarm black hole family) is said to be one of the two conditions at which the black hole resides in a state of thermodynamic equilibrium [22, 26]. In the lukewarm state, the magnitudes of the surface gravities of the event and cosmological horizons are equal. However, equal surface gravities corresponds to equal Hawking temperatures. In the effective temperature formalism, an observer in a lukewarm metric will not notice an equilibrium in the effective temperatures of either the ingoing or the outgoing modes. The cosmological, event and inner horizons are still at distinct spatial radii. If, given that $\varkappa_{c}=-\varkappa_{+}$, the observer were to perceive an equilibrium, one of the requirements would be that
$\frac{\Delta_{ob}}{\Delta_{0}}=\left(\frac{\varkappa_{+}-\varkappa_{ob}}{\varkappa_{+}}\right)\left(\frac{\varkappa_{c}-\varkappa_{ob}}{\varkappa_{c}}\right),$ (30)
with $\varkappa_{ob}$ as defined in (22), and
@@
Figure 2: Density plot of $\log(\kappa^{+}_{i}\Delta)$ as a function of mass and charge in the limit as $r\xrightarrow{}r_{-}$. The plotted region demarcates the realistically permitted values for M and Q. A finite value of $\kappa^{+}\Delta$ as $\Delta\xrightarrow{}0$ implies that $\kappa^{+}$ diverges. A highly negative value of $\kappa^{+}\Delta$ (tending to $\infty$- ∞) indicates a finite effective temperature.
@@
The second region of thermal equilibrium, as stated in literature [26, 27, 28], is the charged Nariai family. It is the extremality determined by the equivalence of $r_{+}$ and $r_{c}$. Furthermore, the surface gravities $\varkappa_{+}=\varkappa_{c}=0$, and the rest point of the observer conincides with these degenerate horizons. With the ingoing and outgoing modes being generated from exactly the same radius, and with the observer having started out from rest at exactly that same point, it is consistent that the rate of redshift/blueshift perceived by the observer is the same for both modes, resulting in the equilibrium. This family demarcates the upper limit on the mass of the black hole, which naturally requires that the event horizon fit within the cosmological boundary of the universe. In an SdS universe, the Nariai branch has no singularities, but with the inclusion of the charge, an effective singularity forms at the inner horizon. The graph on the lower left of Figure 1, with parameters $M=0.25$ and $Q=0.25$, falls at the intersection of the lukewarm and Nariai branches. It shows a state of complete thermodynamic equilibrium, where an observer anywhere in spacetime will see exactly the same effective temperatures if they looked at the sky above, or at the family of emitters below on the event horizon. Note that the equilibrium arises from the Nariai state, rather than from the lukewarm condition. With the event and cosmological horizons coinciding, the rest point of the observer is degenerate with the two horizons, so that that specific energy of the observer is also zero. In such a case, it is not hard to see that this is one of the few cases under which Eq. (30) must always hold true. Following the limits of Eq. (25), the effective temperature in asymptotically dS spacetime tends to 1 (dS length units). It has been calculated, in several papers [29, 25, 22, 27], that the effective temperature of the event and cosmological horizons goes to zero in the Nariai limit. It should be noted here, that the derivation of the “effective temperature” as in the above mentioned papers, stems from a thermodynamic perspective, unlike the present formalism. Regardless, their statement holds true given a standard normalization of the effective temperature. Our formalism inherently developed upon the Bousso-Hawking normalization [30], where the normalization constant ($\gamma_{t}$) in the time-like Killing vector ($\xi=\gamma_{t}\frac{\partial}{\partial t}$) is chosen to satisfy $\xi^{\mu}\xi_{\mu}=-1$ at the equilibrium point of the horizon function (which we defined to be the rest point of our observer) [31].
@@
The range in between the lukewarm and the $r_{+}=r_{-}$ families, the extent of which is dependent on the strength of the cosmological constant, is the “cold” region. It is characterized by a charge marginally greater than the mass, as in lower right graph of Figure 1, given by parameters M=0.25 and Q=0.26. Once again, the transition for the outgoing modes occurs outside the event horizon (which is not unexpected, given that the charge is greater than the mass). The transition for the ingoing modes occurs very close to the inner horizon. The large values of both the mass and the cosmological will contribute a very strong redshift effect, and with the effects of the repulsion from the inner horizon being weaker on the ingoing modes, they remain positive in effective temperature until they get very close to the inner horizon. At the super-extremal limit, where the inner and event horizons coincide exactly, the observer will never see ingoing modes of a negative effective temperature.
@@
Having often used the term “negative effective temperatures”, it is worth taking a moment to consider the implications of a negative effective temperature and therein the importance of the transition points. Starting at the ground state, an increase in energy corresponds to the population of higher energy states, resulting in an increase in entropy, and this process can go on until the system reaches a state of maximum entropy (or infinite temperature). But the energy can be further increased if high energy states are more populated than the low energy states, a phenomenon commonly known as population inversion. At this point, the entropy decreases with an increase in energy, and results in a negative temperature. At maximum entropy, a discontinuity occurs as the temperature jumps from positive to negative infinity [34, 35, 36]. A paper in 1988 [37], perhaps one of the earliest on the topic of negative temperatures in a gravitational context, noted that self gravitating spheres of black body radiation could exhibit alternating regions of positive and negative heat capacities, with the discontinuity at infinite temperature being termed a “phase transition”. Physically, negative temperatures are hotter than positive temperatures. In the context of a black hole, a negative effective temperature means that the observer will no longer perceive Hawking radiation as the Planckian spectrum of a blackbody. The exponential decay of the radiation at high frequencies instead becomes a divergence [38]. In our analysis, a negative effective temperature results from the blueshifting of the Hawking modes, that results from the presence of a horizon of negative surface gravity.
@@ III.3 Divergence at the inner horizon and on the validity of the SCC conjecture
The graphs from Figure 1 suggest a divergence in the outgoing modes at the inner horizon. But these graphs represent only four of an infinite number of possible combinations of M and Q over the entire physically permissible parameter space of the RNdS metric. Furthermore, although the outgoing modes certainly appear to diverge, there is no way, on the basis of the graphs alone, of conclusively determining whether they are diverging to infinity, or tending to a very large finite value. The sharkfin in Figure 2 tests this divergence at the inner horizon for the entire physically permissible parameter space, by plotting a density function of $\log(\kappa^{+}_{i}\Delta)$ over this region.
@@
The converse must also be true (given that the asymptotic form of the horizon function will undergo no significant higher order corrections under extremal conditions)—if the product $\kappa^{+}_{i}\Delta$ is finite, the effective temperature of the outgoing modes must diverge as $(r-r_{-})^{-1}$. This region of infinite divergence is shown in Figure 2, which can be seen to span the majority of the physically permissible parameter space, and all of the non-extremal regimes.
@@
As the surface gravity of the inner horizon approaches zero (which occurs at the extremal cold limit), the second correction term in Eq. (29) starts to become predominant, and the surface gravity of the inner horizon may be corrected as $\varkappa_{-}\approx\varkappa^{\prime}_{-}(r-r_{-})$. The second term scales as $(r-r_{-})^{-1}$, and tends to +$\infty$+ ∞—countering the divergence of the first term to $\infty$- ∞. The effective temperature, in this limit is finite (as can also be seen from Eq. (24a)). The product $\kappa^{+}_{i}\Delta$ tends to zero. This can be seen as the dark line on the extremal cold limit of the sharkfin (close the the $M=Q$ region). It would therefore appear that the SCC conjecture is being violated in the extremal cold regime.
@@
The Nariai limit ($r_{+}=r_{c}$) represented by the right curved edge on the sharkfin, maintains the divergence in the effective temperature. As the charge increases, the magnitudes of both $\varkappa_{+}$ and $\varkappa_{-}$ decreases (and therein the difference between them), the magnitude of the product reduces, and the curve starts to darken. At the ultracold limit ($r_{-}=r_{+}=r_{c}$ at $M=\sqrt{2}/3\sqrt{3}$), the product goes to zero, and the logarithm of the product tends to $\infty$- ∞. In the SdS limit, the surface gravity of the inner horizon approaches $\infty$- ∞, so that the right side of Eq. (34) approaches +$\infty$+ ∞. This can be seen in the sharkfin as the increasingly whiter region towards the lower edge.
@@
Attempts have been made to calculate $\beta$ numerically for quasinormal modes in the permitted RNdS spacetime [49, 42], and regions have been found where it does exceed $1/2$. On this basis, the possibility of a non-divergent behavior was postulated for three conditions: if $M=Q$ [48, 42], in the limit of $\varkappa_{-}=\varkappa_{c}$ [17], or at the extremal cold limit. The effective temperature formalism upholds the SCC conjecture in the lukewarm limit, but fails in the other two. It is able to limit the regions of possible violation to the condition $\varkappa_{+}=\varkappa_{-}=0$, which is a step further than what the calculation of $\beta$ had predicted. Evidently, formulating the effective temperature as a rate of gravitational redshift is able to account for factors that $\beta$ cannot. But a conclusive identification of these factors will perhaps require a more detailed investigation into the relation between $\beta$ and $\kappa$.

### v2

@@ III.2 Effective temperature perceived by a radial freefaller with positive energy
Before proceeding to examine the extremal conditions of the RNdS metric, we may understand the features of the effective temperature in a Schwarzschild de Sitter (SdS) model, which is a simplification (Q=0) of the same metric. It has only the event and cosmological horizons, thereby eliminating the intricacies that accompany the Cauchy horizon. It is therefore a good way to examine the individual effects of the mass and cosmological expansion on the effective temperature. The first graph, on the upper left of Fig. 2, shows the effective temperature for an SdS black hole as a function of the observer’s position. Both the ingoing and outgoing modes are continuous across the event and cosmological horizons, in accordance with the Eqs. (27) and (28). This means that an ingoing(outgoing) observer will be able to cross the event(cosmological) horizon without noticing a discontinuity in the effective temperature. Neither the event nor the cosmological horizons act as a physical barrier, and are simply the radii at which space flows inward (or outward) at the speed of light. As the observer moves to a predominantly de Sitter (dS) space, the effective temperatures of the ingoing and outgoing modes converge to $1/L_{dS}$. In a purely dS space, then, the observer sees an isotropic equilibrium in both the modes, with the value of the effective temperature in this regime being independent of both the (positive) energy and the position of the observer. The energy-independence is in support of the observer-dependence of particle production in dS space [18]. The position-independence reinforces a homogeneity in the dS spacetime. At the other extreme, as the observer approaches the singularity, they will see a monotonic increase, and an asymptotic convergence in the positive effective temperatures of the modes. In the effective temperature formalism, a positive effective temperature corresponds to a perceived positive rate of redshift in the modes. As the curvature of spacetime increases toward the singularity, both the modes are pulled inward, and an infalling observer perceives an increasing rate of redshift. So also, the redshift in the modes in a purely dS spacetime may be understood as resulting from the increasing rate of the expansion of space in this regime. The mass of the black hole and the cosmological expansion, individually, only cause a redshift. The blueshifting, that is seen in the other graphs of Fig. 2, must then be attributed entirely to the effects of the inner horizon.
@@
A very typical representation of an RNdS black hole is seen in the upper right graph of Fig. 2, with parametric values of $M=0.05$, $Q=0.02$. Away from the inner horizon, it is fairly similar to the SdS case. Once again, there is a continuity across the event and cosmological horiozns, and the modes converge to unity in asymptotically dS space. The key feature of this graph, however, comes from the negative effective temperatures of the modes perceived by an observer approaching the inner horizon. A similarity in the cosmological and inner horizons stems from the fact that both have negative surface gravities. But unlike at the cosmological horizon, an observer who started from rest at the maximum of the horizon function and is subject to no external force agents, near the inner horizon is infalling (and ingoing). Effectively, the observer is falling inward and the modes are being repelled by the surface, resulting in a blueshift (which translates to a negative effective temperature). The transition from positive to negative effective temperatures must occur when the redshift caused by the event and cosmological horizons is overcome by the blueshift due to the inner horizon. This in turn, should have a dependency on the closeness of $r_{+}$ and $r_{-}$. Mathematically, the transition occurs at $\varkappa(r_{ob})=\varkappa_{+}$ for the outgoing modes and $\varkappa(r_{ob})=\varkappa_{c}$ for the ingoing modes. While the negative surface gravity certainly causes a blueshift, the effects are much stronger on the outgoing modes than on the ingoing modes because the repulsion supplements their intended direction. As a result, the outgoing modes diverge negatively, while the ingoing modes are blueshifted, but remain finite across the inner horizon.
@@
Having studied the general properties of a very typical RNdS black hole, we may now turn to the more extremal situations. A sharkfin diagram representing the physically permitted parameter space for the metric is plotted in Fig. 3. There are primarily, three regions of interest: the SdS limit (Q=0), the charged Nariai limit ($r_{+}=r_{c}$) and the super-extremal limit ($r_{+}=r_{-}$). In asymptotically flat spacetime, setting M=Q results in an extremal black hole, with a degeneracy in the event and inner horizons. The existence of the cosmological constant raises the upper bounds of the extremality, so that black holes may exist in a limited range where $Q>M$ (the cold region). M=Q (referred to as the lukewarm black hole family) is said to be one of the two conditions at which the black hole resides in a state of thermodynamic equilibrium [24, 29]. In the lukewarm state, the magnitudes of the surface gravities of the event and cosmological horizons are equal. However, equal surface gravities corresponds to equal Hawking temperatures. In the effective temperature formalism, an observer in a lukewarm metric will not notice an equilibrium in the effective temperatures of either the ingoing or the outgoing modes. The cosmological, event and inner horizons are still at distinct spatial radii. If, given that $\varkappa_{c}=-\varkappa_{+}$, the observer were to perceive an equilibrium, one of the requirements would be that
$\frac{\Delta(r_{ob})}{\Delta_{0}}=\left(\frac{\varkappa_{+}-\varkappa(r_{ob})}{\varkappa_{+}}\right)\left(\frac{\varkappa_{c}-\varkappa(r_{ob})}{\varkappa_{c}}\right),$ (32)
with $\varkappa(r_{ob})$ as defined in Eq. (24), and
@@
Figure 3: Density plot of $\log(\kappa^{+}_{i}\Delta)$ as a function of mass and charge in the limit as $r\xrightarrow{}r_{-}$. The plotted region demarcates the realistically permitted values for M and Q. A finite value of $\kappa^{+}\Delta$ as $\Delta\xrightarrow{}0$ implies that $\kappa^{+}$ diverges. A highly negative value of $\kappa^{+}\Delta$ (the log tends to $\infty$- ∞) indicates a finite effective temperature.
@@
The second region of thermal equilibrium, as stated in literature [29, 30, 31], is the charged Nariai family. It is the extremality determined by the equivalence of $r_{+}$ and $r_{c}$. Furthermore, the surface gravities $\varkappa_{+}=\varkappa_{c}=0$, and the rest point of the observer coincides with these degenerate horizons. With the ingoing and outgoing modes being generated from exactly the same radius, and with the observer having started out from rest at exactly that same point, it is consistent that the rate of redshift/blueshift perceived by the observer is the same for both modes, resulting in the equilibrium. This family demarcates the upper limit on the mass of the black hole, which naturally requires that the event horizon fit within the cosmological boundary of the universe. In an SdS universe, the Nariai branch has no singularities, but with the inclusion of the charge, an effective singularity forms at the inner horizon. The graph on the lower left of Fig. 2, with parameters $M=0.25$ and $Q=0.25$, falls at the intersection of the lukewarm and Nariai branches. It shows a state of complete thermodynamic equilibrium, where an observer anywhere in spacetime will see exactly the same effective temperatures if they looked at the sky above, or at the family of emitters below on the event horizon. Note that the equilibrium arises from the Nariai state, rather than from the lukewarm condition. With the event and cosmological horizons coinciding, the rest point of the observer is degenerate with the two horizons, so that the specific energy of the observer is also zero. In such a case, it is not hard to see that this is one of the few cases under which Eq. (32) must always hold true. Following the limits of Eq. (27), the effective temperature in asymptotically dS spacetime tends to 1 (dS length units). It has been calculated, in several papers [32, 28, 24, 30], that the effective temperature of the event and cosmological horizons goes to zero in the Nariai limit. It should be noted here, that the derivation of the “effective temperature” as in the above mentioned papers, stems from a thermodynamic perspective, unlike the present formalism. Regardless, their statement holds true given a standard normalization of the effective temperature. Our formalism inherently developed upon the Bousso-Hawking normalization [33], where the normalization constant ($\gamma_{t}$) in the timelike Killing vector ($\xi=\gamma_{t}\frac{\partial}{\partial t}$) is chosen to satisfy $\xi^{\mu}\xi_{\mu}=-1$ at the equilibrium point of the horizon function (which we defined to be the rest point of our observer) [34].
@@
The range in between the lukewarm and the $r_{+}=r_{-}$ families, the extent of which is dependent on the strength of the cosmological constant, is the “cold” region. It is characterized by a charge marginally greater than the mass, as in lower right graph of Figure 2, given by parameters M=0.25 and Q=0.26. Once again, the transition for the outgoing modes occurs outside the event horizon (which is not unexpected, given that the charge is greater than the mass). The transition for the ingoing modes occurs very close to the inner horizon. The large values of both the mass and the cosmological will contribute a very strong redshift effect, and with the effects of the repulsion from the inner horizon being weaker on the ingoing modes, they remain positive in effective temperature until they get very close to the inner horizon. At the superextremal limit, where the inner and event horizons coincide exactly, the observer will never see ingoing modes of a negative effective temperature.
@@
Having often used the term “negative effective temperatures,” it is worth taking a moment to consider the implications of a negative effective temperature and therein the importance of the transition points. Starting at the ground state, an increase in energy corresponds to the population of higher energy states, resulting in an increase in entropy, and this process can go on until the system reaches a state of maximum entropy (or infinite temperature). But the energy can be further increased if high energy states are more populated than the low energy states, a phenomenon commonly known as population inversion. At this point, the entropy decreases with an increase in energy, and results in a negative temperature. At maximum entropy, a discontinuity occurs as the temperature jumps from positive to negative infinity [37, 38, 39]. A paper in 1988 [40], perhaps one of the earliest on the topic of negative temperatures in a gravitational context, noted that self gravitating spheres of black body radiation could exhibit alternating regions of positive and negative heat capacities, with the discontinuity at infinite temperature being termed a “phase transition”. Physically, negative temperatures are hotter than positive temperatures. In the context of a black hole, a negative effective temperature means that the observer will no longer perceive Hawking radiation as the Planckian spectrum of a blackbody. The exponential decay of the radiation at high frequencies instead becomes a divergence [41]. In our analysis, a negative effective temperature results from the blueshifting of the Hawking modes, that results from the presence of a horizon of negative surface gravity.
@@ III.3 Divergence at the inner horizon and on the validity of the SCC conjecture
The graphs from Fig. 2 suggest a divergence in the outgoing modes at the inner horizon. But these graphs represent only four of an infinite number of possible combinations of M and Q over the entire physically permissible parameter space of the RNdS metric. Furthermore, although the outgoing modes certainly appear to diverge, there is no way, on the basis of the graphs alone, of conclusively determining whether they are diverging to infinity, or tending to a very large finite value. The sharkfin in Fig. 3 tests this divergence at the inner horizon for the entire physically permissible parameter space, by plotting a density function of $\log(\kappa^{+}_{i}\Delta)$ over this region.
@@
The converse must also be true (given that the asymptotic form of the horizon function will undergo no significant higher order corrections under extremal conditions)—if the product $\kappa^{+}_{i}\Delta$ is finite, the effective temperature of the outgoing modes must diverge as $(r-r_{-})^{-1}$. This region of infinite divergence is shown in Figure 3, which can be seen to span the majority of the physically permissible parameter space, and all of the nonextremal regimes.
@@
As the surface gravity of the inner horizon approaches zero (which occurs at the extremal cold limit), the second correction term in Eq. (31) starts to become predominant, and the surface gravity of the inner horizon may be corrected as $\varkappa_{-}\approx\varkappa^{\prime}_{-}(r-r_{-})$. The second term scales as $(r-r_{-})^{-1}$, and tends to +$\infty$+ ∞—countering the divergence of the first term to $\infty$- ∞. The effective temperature, in this limit is finite as can also be seen from Eq. (26a). The product $\kappa^{+}_{i}\Delta$ tends to zero. This can be seen as the dark line on the extremal cold limit of the sharkfin (close the $M=Q$ region). It would therefore appear that the SCC conjecture is being violated in the extremal cold regime.
@@
The Nariai limit ($r_{+}=r_{c}$) represented by the right curved edge on the sharkfin, maintains the divergence in the effective temperature. As the charge increases, the magnitudes of both $\varkappa_{+}$ and $\varkappa_{-}$ decreases (and therein the difference between them), the magnitude of the product reduces, and the curve starts to darken. At the ultracold limit ($r_{-}=r_{+}=r_{c}$ at $M=\sqrt{2}/3\sqrt{3}$), the product goes to zero, and the logarithm of the product tends to $\infty$- ∞. In the SdS limit, the surface gravity of the inner horizon approaches $\infty$- ∞, so that the right side of Eq. (36) approaches +$\infty$+ ∞. This can be seen in the sharkfin as the increasingly whiter region toward the lower edge.
In both the above cases, the limit of extremality has been discussed. Extremal black holes are topologically distinct from nonextremal black holes. One may therefore question the validity in considering the quantum field of an extremal black hole as being a continuity of the nonextremal case. Some papers [54] have clearly claimed that the extremal case in no sense represents a limit of the nonextremal case but implies a real discontinuity. Under certain conditions, a continuity may be maintained as shown in [55], but this may not be generically true. The discussion above should be understood as an (analytical) analysis of subextremal black holes as they approach asymptotically close to the extremal limit. The existence of purely extremal black holes is curtailed by the third law of thermodynamics, and they do not have a realistic channel of formation.
@@
Attempts have been made to calculate $\beta$ numerically for quasinormal modes in the permitted RNdS spacetime [52, 45], and regions have been found where it does exceed $1/2$. On this basis, the possibility of a nondivergent behavior was postulated for three conditions: if $M=Q$ [51, 45], in the limit of $\varkappa_{-}=\varkappa_{c}$ [17], or at the extremal cold limit. The effective temperature formalism upholds the SCC conjecture in the lukewarm limit, but fails in the other two. It is able to limit the regions of possible violation to the condition $\varkappa_{+}=\varkappa_{-}=0$, which is a step further than what the calculation of $\beta$ had predicted. Evidently, formulating the effective temperature as a rate of gravitational redshift is able to account for factors that $\beta$ cannot. But a conclusive identification of these factors will perhaps require a more detailed investigation into the relation between $\beta$ and $\kappa$.